OpenStudy (anonymous):

A farmer has 428 feet of fencing with which he wants to enclose a rectangular garden. What dimensions for the garden will yield a maximum area?

OpenStudy (zzr0ck3r):

area = x*y paremeter = 2x+2y= 428 x = 224-y area = x*y = (224-y)*y area = -y^2 + 224y do you know how to maxemise that?

OpenStudy (anonymous):

-2y+224?

OpenStudy (anonymous):

i think that's 214 not 224....

OpenStudy (zzr0ck3r):

yeah 214:)

OpenStudy (zzr0ck3r):

@jeanniebean do you know how to find a virtex?

OpenStudy (anonymous):

no

OpenStudy (zzr0ck3r):

we know that it "opens down" so it will have some maximum point at the "top" of the paralbola

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):



OpenStudy (zzr0ck3r):

there is a formula for that its -b/2a where b and a are the coefficients on the 1 and second term respectivly

OpenStudy (zzr0ck3r):

-214/(-2*1) = 214/2 = 107

OpenStudy (anonymous):

thats what i got but i dont think i went the same way of getting it

OpenStudy (zzr0ck3r):

so x = 107 use this to find y and x*y = the answer

OpenStudy (zzr0ck3r):

there are many ways to do it:)

OpenStudy (anonymous):

isnt it 107x107

OpenStudy (zzr0ck3r):

no

OpenStudy (anonymous):

darnit

OpenStudy (zzr0ck3r):

wait yes

OpenStudy (anonymous):

yes it is... the maximum area for a rectangle with the given perimeter is a square.

OpenStudy (anonymous):

ok cool!

OpenStudy (zzr0ck3r):

sorry pizza just got here I'm doing to many things at once

OpenStudy (anonymous):

mmm... pizza... can i have some?

OpenStudy (zzr0ck3r):

if your in ne portland:)

OpenStudy (anonymous):

ok thanks! i have a few more that i need help with so i will post those too! and thats what im saying all this math is making me starving!

OpenStudy (zzr0ck3r):

ha:)